# Exactly solvable Schrödinger operators related to the hypergeometric equation

### v1

@@ 5.2 Hyperbolic case
$\mathbb{P}_{\alpha,\beta,\mu}^{\rm h}(u)$ $:=({\rm e}^{2u}-1)^{-\frac{\alpha}{2}}(1-{\rm e}^{-2u})^{-\frac{\beta}{2}}~\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha+\beta-\mu+1}{2};\alpha+1;\frac{1}{1-{\rm e}^{2u}}\right)$ (5.47)
$=({\rm e}^{2u}-1)^{-\frac{\alpha}{2}}(1-{\rm e}^{-2u})^{\frac{1+\alpha+\mu}{2}}~\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha-\beta+\mu+1}{2};\alpha+1;{\rm e}^{-2u}\right),$ (5.48)
$\mathbb{Q}_{\alpha,\beta,\mu}^{\rm h}(u)$ $:=({\rm e}^{2u}-1)^{\frac{1+\beta+\mu}{2}}(1-{\rm e}^{-2u})^{-\frac{\beta}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{-\alpha+\beta+\mu+1}{2};\mu+1;1-{\rm e}^{2u}\right)$ (5.49)
$=({\rm e}^{2u}-1)^{-\frac{\alpha}{2}}(1-{\rm e}^{-2u})^{\frac{1+\alpha+\mu}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha-\beta+\mu+1}{2};\mu+1;1-{\rm e}^{-2u}\right).$ (5.50)
@@
$\sigma_{\rm d}(K_{\kappa,\mu}^{\rm h})=$ $\bigg{\{}-\frac{\kappa^{2}}{(2k+\mu)^{2}}-\Big{(}k+\frac{\mu}{2}\Big{)}^{2}\quad\bigg{|}\quad k\in{\mathbb{N}}_{0}+\frac{1}{2},\quad{\rm Re}\left(\frac{\kappa}{2k+\mu}+k+\frac{\mu}{2}\right)<0\bigg{\}},$
@@
${\cal W}\left(\mathbb{Q}^{\rm h}_{\alpha,\beta,\mu},\mathbb{P}^{\rm h}_{\alpha,\beta,\mu}\right)$ $=2{\cal W}\left(Q^{\rm h}_{\alpha,\beta,\mu},P^{\rm h}_{\alpha,\beta,\mu}\right)$ (5.63)
@@
For $\mu\in{\mathbb{Z}}$ we have an additional identity for the $\mathbb{Q}^{\rm h}$ function, which follows directly from (A.17):
@@
$\frac{1}{(K_{\tau,1}^{\rm h}+\delta)}=\frac{1}{(K_{\tau,-1}^{\rm h}+\delta)}.$ (5.68)
@@ 5.3 DeSitterian case
$\mathbb{P}_{\alpha,\beta,\mu}^{\mathrm{dS}}(u)$ $:=(1+{\rm e}^{2u})^{-\frac{\alpha}{2}}(1+{\rm e}^{-2u})^{-\frac{\beta}{2}}\mathbf{F}\Big{(}\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha+\beta-\mu+1}{2};\alpha+1;\frac{1}{1+{\rm e}^{2u}}\Big{)}$ (5.74)
$:=(1+{\rm e}^{2u})^{-\frac{\alpha}{2}}(1+{\rm e}^{-2u})^{\frac{1+\alpha+\mu}{2}}\mathbf{F}\Big{(}\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha-\beta+\mu+1}{2};\alpha+1;-{\rm e}^{-2u}\Big{)}.$ (5.75)
@@
$\frac{1}{(K_{\kappa,\mu}^{\mathrm{dS}}+\delta)}(x,y)=$ $\frac{\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta-\mu}{2}\right)}{2}$
@@
${\cal W}\left(\mathbb{P}^{\mathrm{dS}}_{\alpha,\beta,\mu}(u),\mathbb{P}^{\mathrm{dS}}_{\beta,\alpha,\mu}(-u)\right)$ $=2{\cal W}\left(P^{{\rm s}}_{\alpha,\beta,\mu}(r),P^{{\rm s}}_{\beta,\alpha,\mu}(\pi-r)\right)$ (5.86)
@@
$\mathbb{P}_{\alpha,\beta,\mu}^{\mathrm{dS}}(w)$ $=\big{(}1-{\rm e}^{-2u}\big{)}^{-\frac{1}{2}}\mathbb{P}_{\alpha,\mu,\beta}^{\rm h}(u),$ (5.91)
$\mathbb{P}_{\beta,\alpha,\mu}^{\mathrm{dS}}(-w)$ $=\big{(}1-{\rm e}^{-2u}\big{)}^{-\frac{1}{2}}\mathbb{Q}_{\alpha,\mu,\beta}^{\rm h}(u).$ (5.92)
@@ 6 The Laplacian on an interval, halfline and line
The Laplacians on $0,\pi[$, $\mathbb{R}$$_{+}$ and $\mathbb{R}$} with the Dirichlet or Neumann boundary conditions at endpoints belong to the most widely used operators. Their Green functions can be easily computed in terms of elementary functions, without using hypergeometric functions. In this section we will check that they are special cases of hypergeometric Hamiltonians. We will see that this coincidence is related to various identities for hypergeometric functions from Appendix A.6.
@@ 6.1 Laplacian on an interval
Let us check that (6.12)–(6.15) agree with the more general formula (4.20) involving the hypergeometric function and the Gamma function. We identify $k=\frac{{\rm i}\mu}{2}$. By (A.32) and (A.33) we obtain
@@ 6.2 Laplacian on the halfline
$(L_{{\mathrm{D}}}+k^{2})^{-1}(x,y)=\frac{1}{k}\begin{cases}\sinh kx~{\rm e}^{-ky},\quad\text{if}\quad 0<x<y,\\ \sinh ky~{\rm e}^{-kx},\quad\text{if}\quad 0<y<x;\end{cases}$ (6.29)
$(L_{{\mathrm{N}}}+k^{2})^{-1}(x,y)=\frac{1}{k}\begin{cases}\cosh kx~{\rm e}^{-ky},\quad\text{if}\quad 0<x<y,\\ \cosh ky~{\rm e}^{-kx},\quad\text{if}\quad 0<y<x.\end{cases}$ (6.30)
To check that (6.29) and (6.30) agree with (4.40), identify $k=\frac{\mu}{2}$. By (A.34), (A.35) and (A.36), we have
@@ 6.3 Laplacian on the line
$(L+k^{2})^{-1}(x,y)=\frac{1}{2k}\begin{cases}{\rm e}^{kx}{\rm e}^{-ky},\quad\text{if}\quad x<y,\\ {\rm e}^{ky}{\rm e}^{-kx},\quad\text{if}\quad y<x.\end{cases}$ (6.41)
@@
which follows from (A.36), and
@@ 7.1 Sphere
$\cos r=w,\quad\sin r=\sqrt{1-w^{2}}.$ (7.6)
In these coordinates we first write the line element, then the Laplacian:
@@ 7.2 Hyperbolic space
$X_{0}=\cosh r,\quad i=1,\dots,d,\quad X_{i}=\sinh r\,\hat{X}_{i}.$ (7.13)
@@
$\cosh r=w,\quad\sinh r=\sqrt{w^{2}-1}.$ (7.14)
In these coordinates we first write the line element, then the Laplacian:
@@ 7.3 DeSitter space
Alternative coordinates are
$\sinh t=w,$ $\quad\cosh t=\sqrt{1+w^{2}}.$ (7.22)
In these coordinates we first write the line element, then the d’Alembertian:
${\rm d}s^{2}$ $=-{\rm d}t^{2}+\cosh^{2}t{\rm d}\hat{X}^{2}$ (7.23)
@@
$\Box_{d}^{\mathrm{dS}}$ $=-\partial_{t}^{2}+(d-1)\tanh t\partial_{t}+\frac{\Delta_{d-1}^{\rm s}}{\cosh^{2}t}$ (7.25)
@@ 7.4 Sphere in double spherical coordinates
$\Delta_{p+q-1}^{\rm s}$ $=4w(1-w)\partial_{w}^{2}+2\big{(}p(1-w)-qw\big{)}\partial_{w}+\frac{\Delta_{p-1}^{\rm s}}{w}+\frac{\Delta_{q-1}^{\rm s}}{1-w}$ (7.34)
@@
$(\sin\tau)^{\frac{p-1}{2}}(\cos\tau)^{\frac{q-1}{2}}\big{(}-\Delta_{p+q-1}^{\rm s}\big{)}(\sin\tau)^{-\frac{p-1}{2}}(\cos\tau)^{-\frac{q-1}{2}}+\big{(}\tfrac{p+q-2}{2}\big{)}^{2}$ (7.36)
@@ 7.5 Hyperboloid in double spherical coordinates
$\Delta_{p-1,q}$ $=4w(w-1)\partial_{w}^{2}+2\big{(}p(w-1)+qw\big{)}\partial_{w}-\frac{\Delta_{p-1}^{\rm s}}{w}+\frac{\Delta_{q-1}^{\rm s}}{w-1}$ (7.45)
@@
$(\cosh\tau)^{\frac{p-1}{2}}(\sinh\tau)^{\frac{q-1}{2}}\big{(}-\Delta_{p-1,q}\big{)}(\cosh\tau)^{-\frac{p-1}{2}}(\sinh\tau)^{-\frac{q-1}{2}}-\big{(}\tfrac{p+q-2}{2}\big{)}^{2}$ (7.47)
@@ 7.6 Complex manifolds
(Note that there is no absolute value).
@@
we have the direct analog of the formula from the real case:
@@
$\Delta_{d,{\mathbb{C}}}^{\rm s}$ $=\sum_{0\leq i<j\leq d}(Z_{i}\partial_{Z_{j}}-Z_{j}\partial_{Z_{i}})^{2}$ (7.62)
@@ 7.7 Hyperboloid ${\mathbb{H}}^{p-1,p}$ in complex coordinates
${\mathbb{H}}^{p-1,p}$ $=\{(X,Y)\in{\mathbb{R}}^{2p}\ |\ -Y_{1}^{2}-\cdots-Y_{p}^{2}+X_{1}^{2}+\cdots+X_{p}^{2}=-1\}$ (7.64)
@@
$-{\rm d}Y_{1}^{2}-\cdots-{\rm d}Y_{p}^{2}+{\rm d}X_{1}^{2}+\cdots+{\rm d}X_{p}^{2}$ (7.66)
$=$ ${\rm d}Z_{1}^{2}+\cdots+{\rm d}Z_{p}^{2}+{\rm d}\overline{Z}_{1}^{2}+\cdots+{\rm d}\overline{Z}_{1}^{2}$ (7.67)
@@
$\big{(}\cosh r)^{\frac{p-1}{2}}\big{(}\Delta_{p-1,p}\big{)}(\cosh r)^{-\frac{p-1}{2}}-(p-1)^{2}$ (7.72)

### v2

@@ 5.2 Hyperbolic case
$\mathbb{P}_{\alpha,\beta,\mu}^{\rm h}(u)$ $:=\left({\rm e}^{2u}-1\right)^{-\frac{\alpha}{2}}\left(1-{\rm e}^{-2u}\right)^{-\frac{\beta}{2}}~\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha+\beta-\mu+1}{2};\alpha+1;\frac{1}{1-{\rm e}^{2u}}\right)$
$=\left({\rm e}^{2u}-1\right)^{-\frac{\alpha}{2}}(1-{\rm e}^{-2u})^{\frac{1+\alpha+\mu}{2}}~\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha-\beta+\mu+1}{2};\alpha+1;{\rm e}^{-2u}\right),$ (5.43)
$\mathbb{Q}_{\alpha,\beta,\mu}^{\rm h}(u)$ $:=\left({\rm e}^{2u}-1\right)^{\frac{1+\beta+\mu}{2}}\left(1-{\rm e}^{-2u}\right)^{-\frac{\beta}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{-\alpha+\beta+\mu+1}{2};\mu+1;1-{\rm e}^{2u}\right)$
$=\left({\rm e}^{2u}-1\right)^{-\frac{\alpha}{2}}\left(1-{\rm e}^{-2u}\right)^{\frac{1+\alpha+\mu}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha-\beta+\mu+1}{2};\mu+1;1-{\rm e}^{-2u}\right).$ (5.44)
@@
$\sigma_{\rm d}(K_{\kappa,\mu}^{\rm h})=$ $\bigg\{-\frac{\kappa^{2}}{(2k+\mu)^{2}}-\Big(k+\frac{\mu}{2}\Big)^{2}\quad\bigg|\quad k\in{\mathbb{N}}_{0}+\frac{1}{2},\quad{\rm Re}\left(\frac{\kappa}{2k+\mu}+k+\frac{\mu}{2}\right)<0\bigg\},$ (5.52)
@@
${\cal W}\left(\mathbb{Q}^{\rm h}_{\alpha,\beta,\mu},\mathbb{P}^{\rm h}_{\alpha,\beta,\mu}\right)$ $=2{\cal W}\left(Q^{\rm h}_{\alpha,\beta,\mu},P^{\rm h}_{\alpha,\beta,\mu}\right)$
@@
For $\mu\in{\mathbb{Z}}$ we have an additional identity for the $\mathbb{Q}^{\rm h}$ function, which follows directly from (C.17):
@@
$\frac{1}{\left(K_{\tau,1}^{\rm h}+\delta\right)}=\frac{1}{\left(K_{\tau,-1}^{\rm h}+\delta\right)}.$ (5.62)
@@ 5.3 DeSitterian case
$\mathbb{P}_{\alpha,\beta,\mu}^{\mathrm{dS}}(u)$ $:=\left(1+{\rm e}^{2u}\right)^{-\frac{\alpha}{2}}\left(1+{\rm e}^{-2u}\right)^{-\frac{\beta}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha+\beta-\mu+1}{2};\alpha+1;\frac{1}{1+{\rm e}^{2u}}\right)$
$=\left(1+{\rm e}^{2u}\right)^{-\frac{\alpha}{2}}\left(1+{\rm e}^{-2u}\right)^{\frac{1+\alpha+\mu}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha-\beta+\mu+1}{2};\alpha+1;-{\rm e}^{-2u}\right).$ (5.68)
@@
$\frac{1}{\left(K_{\kappa,\mu}^{\mathrm{dS}}+\delta\right)}(x,y)=$ $\frac{\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta-\mu}{2}\right)}{2}$
@@
${\cal W}\left(\mathbb{P}^{\mathrm{dS}}_{\alpha,\beta,\mu}(u),\mathbb{P}^{\mathrm{dS}}_{\beta,\alpha,\mu}(-u)\right)$ $=2{\cal W}\left(P^{{\rm s}}_{\alpha,\beta,\mu}(r),P^{{\rm s}}_{\beta,\alpha,\mu}(\pi-r)\right)$
@@
$\mathbb{P}_{\alpha,\beta,\mu}^{\mathrm{dS}}(w)$ $=\left(1-{\rm e}^{-2u}\right)^{-\frac{1}{2}}\mathbb{P}_{\alpha,\mu,\beta}^{\rm h}(u),$ (5.83)
$\mathbb{P}_{\beta,\alpha,\mu}^{\mathrm{dS}}(-w)$ $=\left(1-{\rm e}^{-2u}\right)^{-\frac{1}{2}}\mathbb{Q}_{\alpha,\mu,\beta}^{\rm h}(u).$ (5.84)
@@ 6 The Laplacian on an interval, halfline and line
The Laplacians on $0,\pi[$, $\mathbb{R}$$_{+}$ and $\mathbb{R}$} with the Dirichlet or Neumann boundary conditions at endpoints belong to the most widely used operators. Their Green functions can be easily computed in terms of elementary functions, without using hypergeometric functions. In this section we will check that they are special cases of hypergeometric Hamiltonians. We will see that this coincidence is related to various identities for hypergeometric functions from Appendix C.6.
@@ 6.1 Laplacian on an interval
Let us check that (6.12)–(6.15) agree with the more general formula (4.19) involving the hypergeometric function and the Gamma function. We identify $k=\frac{{\rm i}\mu}{2}$. By (C.31) and (C.32) we obtain
@@ 6.2 Laplacian on the halfline
$\left(L_{{\mathrm{D}}}+k^{2}\right)^{-1}(x,y)=\frac{1}{k}\begin{cases}\sinh kx~{\rm e}^{-ky},\quad\text{if}\quad 0<x<y,\\ \sinh ky~{\rm e}^{-kx},\quad\text{if}\quad 0<y<x;\end{cases}$ (6.29)
$\left(L_{{\mathrm{N}}}+k^{2}\right)^{-1}(x,y)=\frac{1}{k}\begin{cases}\cosh kx~{\rm e}^{-ky},\quad\text{if}\quad 0<x<y,\\ \cosh ky~{\rm e}^{-kx},\quad\text{if}\quad 0<y<x.\end{cases}$ (6.30)
To check that (6.29) and (6.30) agree with (4.36), identify $k=\frac{\mu}{2}$. By (C.33), (C.34) and (C.35), we have
@@ 6.3 Laplacian on the line
$\left(L+k^{2}\right)^{-1}(x,y)=\frac{1}{2k}\begin{cases}{\rm e}^{kx}{\rm e}^{-ky},\quad\text{if}\quad x<y,\\ {\rm e}^{ky}{\rm e}^{-kx},\quad\text{if}\quad y<x.\end{cases}$ (6.41)
@@
which follows from (C.35), and
@@ 7.2 Hyperbolic space
$X_{0}=\cosh r,\quad i=1,\dots,d,\quad X_{i}=\sinh r\,\hat{X}_{i},$ (7.12)
@@ 7.3 DeSitter space
${\rm d}s^{2}$ $=-{\rm d}t^{2}+\cosh^{2}t{\rm d}\hat{X}^{2}$
@@
$\Box_{d}^{\mathrm{dS}}$ $=-\partial_{t}^{2}+(d-1)\tanh t\partial_{t}+\frac{\Delta_{d-1}^{\rm s}}{\cosh^{2}t}$
@@ 7.4 Sphere in double spherical coordinates
$\Delta_{p+q-1}^{\rm s}$ $=4w(1-w)\partial_{w}^{2}+2\big(p(1-w)-qw\big)\partial_{w}+\frac{\Delta_{p-1}^{\rm s}}{w}+\frac{\Delta_{q-1}^{\rm s}}{1-w}$
@@
$(\sin\tau)^{\frac{p-1}{2}}(\cos\tau)^{\frac{q-1}{2}}\big(-\Delta_{p+q-1}^{\rm s}\big)(\sin\tau)^{-\frac{p-1}{2}}(\cos\tau)^{-\frac{q-1}{2}}+\big(\tfrac{p+q-2}{2}\big)^{2}$
@@ 7.5 Hyperboloid in double spherical coordinates
$\Delta_{p-1,q}$ $=4w(w-1)\partial_{w}^{2}+2\big(p(w-1)+qw\big)\partial_{w}-\frac{\Delta_{p-1}^{\rm s}}{w}+\frac{\Delta_{q-1}^{\rm s}}{w-1}$
@@
$(\cosh\tau)^{\frac{p-1}{2}}(\sinh\tau)^{\frac{q-1}{2}}\big(-\Delta_{p-1,q}\big)(\cosh\tau)^{-\frac{p-1}{2}}(\sinh\tau)^{-\frac{q-1}{2}}-\big(\tfrac{p+q-2}{2}\big)^{2}$
@@ 7.6 Complex manifolds
Note that there is no absolute value, and $\det g$ is always non-zero. The definition of $\Delta$ does not depend on the choice of the branch of the (double-valued) square root.
@@
where we arbitrarily select the branch of the square root, we have the direct analog of the formula from the real case:
@@
$\Delta_{d,{\mathbb{C}}}^{\rm s}$ $=\sum_{0\leq i<j\leq d}(Z_{i}\partial_{Z_{j}}-Z_{j}\partial_{Z_{i}})^{2}$
@@ 7.7 Hyperboloid ${\mathbb{H}}^{p-1,p}$ in complex coordinates
${\mathbb{H}}^{p-1,p}$ $=\{(X,Y)\in{\mathbb{R}}^{2p}\ |\ -Y_{1}^{2}-\cdots-Y_{p}^{2}+X_{1}^{2}+\cdots+X_{p}^{2}=-1\}$
@@
$-{\rm d}Y_{1}^{2}-\cdots-{\rm d}Y_{p}^{2}+{\rm d}X_{1}^{2}+\cdots+{\rm d}X_{p}^{2}$
$=$ ${\rm d}Z_{1}^{2}+\cdots+{\rm d}Z_{p}^{2}+{\rm d}\overline{Z}_{1}^{2}+\cdots+{\rm d}\overline{Z}_{1}^{2}$
@@
$\big(\cosh r)^{\frac{p-1}{2}}\big(\Delta_{p-1,p}\big)(\cosh r)^{-\frac{p-1}{2}}-(p-1)^{2}$